OpenStudy (anonymous):

solve using the multiplication principle -18=19y Im lost with this one

OpenStudy (anonymous):

the multiplication rule explains that if you multiply or divide both sides of an equatio nby the same number you arrive at the same answer. so, 5x = 3y is the same as (5x)/5 = (3y)/5

OpenStudy (anonymous):

Inf your case, -18 = 19y, but ALSO, -18/19 = 19y/19 this is helpful because 19y/19 equals y, so now you've solved for y by not changing the equation illegally

OpenStudy (anonymous):

so "solving" an equation just means multiplying/dividing by whatever you need to on both side of the equation using a number that will get rid of another number. in this case, dividing both sides by 19 leaves y all by itself.

OpenStudy (anonymous):

ok so my answer would be 19

OpenStudy (anonymous):

well nope, not quite.

OpenStudy (anonymous):

you're dividing each side by 19 but y = -18/19

OpenStudy (anonymous):

so -18/19 is your answer. that is what "solving" using the multiplication principle means here.

OpenStudy (anonymous):

so because 19y/19 equals y this is why the letter y isn't in the final answer?